Circular No. 7611

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ay IN IC 4423 B. Swift and W. D. Li, University of California at Berkeley, report the

discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 16.9) on an unfiltered CCD image taken on Apr. 18.4 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001ay, which was confirmed at mag ∼ 17.0 on an image taken on Apr. 19.2 by M. Schwartz with the 0.5-m Tenagra III automatic telescope, is located at α = 14^{h}26^{m}17^{s}.00, δ = +26^{o}14′55′′.8 (equinox 2000.0), which is 10′′.3 west and 9′′.3 north of the nucleus of IC 4423. A KAIT image of the same field taken on Apr. 4.4 showed nothing at this position (limiting mag ∼ 18.5).

SGR 1900+14 C. Guidorzi, F. Frontera, and E. Montanari, University of Ferrara;

L. Amati and M. Orlandini, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; E. Costa, M. Feroci, and G. Gandolfi, Istituto Astrofisica Spaziale, CNR, Rome; J. J. M. in ’t Zand, Astronomical Institute, University of Utrecht and Space Research Organization Netherlands; and G. D’Andreta, BeppoSAX Science Operation Center, report: “A giant flare from SGR 1900+14 was detected by BeppoSAX on Apr. 18.33000 UT. In the Gamma Ray Burst Monitor (GRBM) energy band (40–700 keV), the flare shows a fast rise (∼ 8 s) and a slow decay, with a total duration of ∼ 40 s. Superposed to the time profile, the 5-s pulsation is apparent. The 40–700keV peak flux, after deadtime correction, is 16 400 counts/s, corresponding to (1.1 ± 0.1) × 10^{−5} erg cm^{−2} s^{−1}. A preliminary analysis shows that the average GRBM spectrum is consistent with a thermal bremsstrahlung (kT ∼ 30 keV) plus a hard tail. Due to the brightness of the flare, the Wide Field Camera unit 1 was automatically switched off after a few seconds. Before that, it showed a peak flux in excess of 17 Crab units (2–28 keV). The source position determined with the WFC is consistent with that of SGR 1900+14, with an accuracy of 2′.1 (99-percent confidence).”

COMET C/2001 A2 (LINEAR) Visual m_{1} estimates: Mar. 31.42 UT, 6.9 (D. A. J. Seargent, The En-

trance, N.S.W., 10×50 binoculars); Apr. 11.76, 7.4 (A. Baransky, Kiev, Ukraine, 0.11-m reflector); 13.84, 7.6 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 14.38, 7.4 (Seargent); 16.16, 7.2 (C. S. Morris, Fillmore, CA, 20×80 binoculars); 18.37, 7.0 (Seargent, 2×20 binoculars).

2001 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green